# ResKoopNet: Learning Koopman Representations for Complex Dynamics with Spectral Residuals

### v3

@@ Abstract
Analyzing long-term behaviors in high-dimensional nonlinear dynamical systems remains challenging, with the Koopman operator framework providing a powerful global linearization approach, though existing methods for approximating its spectral components often suffer from theoretical limitations and reliance on predefined dictionaries. While Residual Dynamic Mode Decomposition (ResDMD) introduced the spectral residual to assess the accuracy of Koopman operator approximation, it only filters precomputed spectra, which prevents it from fully discovering the Koopman operator’s complete spectral information (a limitation sometimes referred to as the ‘spectral inclusion’ problem). We introduce ResKoopNet (Residual-based Koopman-learning Network), a novel method that addresses this limitation by explicitly minimizing the spectral residual to compute Koopman eigenpairs, which can identify a more precise and complete spectrum of the Koopman operator. This approach provides theoretical guarantees while maintaining computational adaptability through a neural network implementation. Experiments on physical and biological systems demonstrate ResKoopNet’s superior accuracy in spectral approximation compared to existing methods, particularly for systems with continuous spectra and high dimensional, which makes it as an effective tool for analyzing complex dynamical systems.
@@ 1 Introduction
To address these limitation, Residual Dynamic Mode Decomposition (ResDMD) [8] was introduced to offer convergence guarantees through a spectral residual that measures how well the estimated Koopman spectra are. By assessing convergence of spectral residual, ResDMD eliminates spurious spectral components that do not represent the system’s true dynamics, which enhances the reliability of spectral estimation. ResDMD also introduces the concept of pseudospectrum, regions where the spectral content is significant even without discrete eigenvalues, which is particularly valuable for analyzing systems with continuous spectra. However, ResDMD only filters precomputed polluted spectra, which limits its ability to independently refine spectral estimates.
@@ 3.1 ResDMD Review
$\begin{split}\widehat{\textit{res}}(\lambda,\phi)^{2}\coloneqq\frac{1}{m}\mathbf{v}^{*}[\Psi_{Y}^{*}\Psi_{Y}-\lambda(\Psi_{X}^{*}\Psi_{Y})^{*}\\ -\bar{\lambda}\Psi_{X}^{*}\Psi_{Y}+|\lambda|^{2}\Psi_{X}^{*}\Psi_{X}]\mathbf{v}.\end{split}$ (3.3)
@@ 3.2 Residual-based Koopman-learning Network (ResKoopNet)
General framework In this section, we present the Residual-based Koopman-learning Network (ResKoopNet), which minimizes spectral residual to address the spectral inclusion problem and yield a more accurate and complete Koopman spectrum. The overall pipeline is illustrated in Figure 1. ResKoopNet optimizes the dictionary functions by minimizing the total spectral residual,
@@
Figure 1: The optimizer on the right looks for the optimal dictionary by minimizing Eq. (3.3).
@@ 4 Application in physical and biological systems
In this section, we present three examples to demonstrate ResKoopNet’s effectiveness in estimating key components of Koopman Mode Decomposition: spectrum, eigenfunctions, and Koopman modes. In the pendulum example, we compare with EDMD, EDMD-DL, Hankel-DMD, and ResDMD, and show that ResKoopNet captures the full Koopman spectrum with significantly fewer dictionary observables than required by ResDMD, while other methods either produce spectral pollution or miss the continuous spectrum. Our turbulence analysis demonstrates ResKoopNet’s ability to detect not only acoustic vibration and turbulent fluctuation (as in ResDMD [8, Section 4.3.1, Section 6.3]) but also to recover the fundamental pressure field structure through the lowest-residual Koopman mode, which is a feature not captured by Hankel-DMD using the same dataset. In the neural dynamics example, we compare against three methods across five mice datasets, with ResKoopNet achieving superior clustering of neural states corresponding to different visual stimuli, as quantified by the Davies-Bouldin index. Note that we include Hankel-DMD [1] as a consistent benchmark across all experiments due to its theoretical guarantees and applicability to both low and high-dimensional systems without requiring predefined dictionaries. It uses time-delayed state measurements by Takens Embedding Theorem [36] instead of using a pre-defined dictionary or applying Galerkin approximation as in EDMD and ResDMD methods (see Appendix A.8.1 for more details.)
@@ 4.1 Pendulum
As shown in Figure 2, ResKoopNet requires only $N_{K}=300$ observables to approximate the full spectrum. While the original ResDMD work reported using 964 basis functions [8, Section 4.3.1], our additional tests indicate that approximately 460 basis functions can produce reasonable circular approximations with ResDMD (see Figure 7 in Appendix A.7.1). Nevertheless, ResKoopNet still achieves comparable accuracy with significantly fewer observables. This efficiency persists even with larger sizes of data points (Figure 3), where $N_{K}=350$ observables remain sufficient. Figure 4 compares ResKoopNet with four classical methods using the 90-initial-point dataset. Both EDMD and ResDMD employ Hermite functions (order 15) for $\theta$ and Fourier functions (order 20) for $\dot{\theta}$, while Hankel-DMD uses a time delay of 150. Though Hankel-DMD produces eigenvalues near the unit circle, it only captures the discrete point spectrum and cannot reconstruct the continuous spectral components. We include Hankel-DMD here for consistency in benchmarking across all experiments, despite its limitations in higher-dimensional systems examined later. The shaded region in the ResDMD plot represents the $\epsilon$-pseudospectrum, which approximates regions where the spectral residual is smaller than a threshold value; however, with 300 basis functions, ResDMD fails to fully recover the unit circle. Overall, this pendulum example demonstrates that ResKoopNet achieves accurate and complete spectral recovery using much fewer observables than other classical methods.
@@ 4.2 Turbulence
Recovering spatial patterns is a typical goal of DMD-based methods, especially in fluid dynamics [33, 28]. Figure 5(a) shows the ground truth pressure distribution, where a clear asymmetry exists between the upper and lower airfoil surfaces. This fundamental spatial pattern is critical for analyzing aerodynamic performance and structural integrity. While the original Kernel-ResDMD method has successfully detected acoustic vibration and turbulent fluctuation shown in Colbrook and Townsend [8, Section 6.3], it fails to recover the global spatial structure of the pressure field. In contrast, ResKoopNet achieves a significant improvement by accurately reconstructing this dominant spatial pressure field structure itself from its leading Koopman mode as shown in Figure 5(b), which contains essential information about the flow characteristics around the airfoil.
@@ 4.3 Neural dynamics identification in mice visual cortex
Since ResKoopNet directly minimizes the residuals based on eigenfunctions, its estimated evolution of eigenfunctions over time should ideally capture latent dynamics. To evaluate how effectively ResKoopNet reveals latent temporal dynamics in real data, we apply it to a dataset of high-dimensional neural signals and demonstrate its advantages over a series of classical methods: the Hankel-DMD, EDMD (combined with RBF basis) and Kernel-ResDMD. These methods are selected as representative approaches for handling high-dimensional data.
@@
We applied ResKoopNet and three classical Koopman decomposition methods (Hankel-DMD, EDMD with RBF basis, and Kernel-ResDMD) to this dataset, using different implementations and Koopman subspace dimensions. For ResKoopNet, we trained dictionaries on all snapshots from each mouse to avoid overfitting. We reduced the data to 24 dimensions via truncated SVD, together with the constant and 25 trainable bases, resulting in 50 eigenfunctions. The neural network we use here consists of 3 hidden layers, each with 200 neurons. The hyperparameter scanning results are included in Appendix Figure 17, demonstrating that our parameter choice is within the robust range. The decomposed eigenfunctions are shown in Figure 6A(top), with markers indicating ground truth state separations. For Hankel-DMD, we built a Hankel matrix with a delay of 50, producing 50 eigenfunctions per trial. In EDMD with RBF basis, we used the SVD-truncated 300 basis and 1000 RBF functions, resulting in 1301 eigenfunctions. For Kernel-ResDMD, we used normalized Gaussians as kernel functions, setting the Koopman subspace dimension to 299 eigenfunctions based on Colbrook et al. [9]. See Appendix A.9.4 for method details and Appendix A.10 for dictionary size justification. These eigenfunctions, shown in Figure 6A(bottom), Appendix Figure 14A, and Appendix Figure 15A, are compared to ground truth trial identities.
The Koopman eigenfunctions represent dynamical features corresponding to the video stimuli. To evaluate their effectiveness, we assess how well eigenfunctions of the same stimuli cluster together, which distinguishes them from other states. If the eigenfunctions capture key dynamics related to the stimuli, those from trials with the same video should be separable from others. This turns the problem into a clustering task based on the separability of eigenfunctions across different stimuli. Note that averaged trial differences are even visibly clear for the ResKoopNet case. It is worth clarifying that clustering these processes is not biologically necessary, as simple clustering of mean firing rates is sufficient (see Appendix Figures 1112). However, this dataset provides an ideal benchmark with clear ground truth labels, enabling accurate comparison of Koopman eigenfunction estimation methods. Effective estimation of Koopman eigenfunctions can be valuable for more complex tasks, such as unsupervised latent state identification or decoding object dynamics from video stimuli, offering promising directions for future research.
@@ 5 Conclusion and future work
In this paper, we introduced ResKoopNet, a novel approach that minimizes spectral residuals to accurately approximate Koopman operators for complex dynamical systems. This method successfully resolves the spectral inclusion problem while capturing both discrete and continuous spectra with high fidelity. Our experiments across physical and biological systems demonstrate that ResKoopNet significantly outperforms existing methods in identifying coherent structures and uncovering latent dynamics, particularly in high-dimensional applications.
Despite these strengths, ResKoopNet has several limitations. First, the use of neural networks increases computational cost compared to classical algorithms (see Appendix A.6). Second, unlike stochastic approaches such as VAMP [25] and SDMD [46], ResKoopNet currently does not account for stochasticity (see Appendix A.5 for comparison). In addition, its performance can vary depending on the neural architecture and training configuration.
Many physical laws, such as conservation principles or symmetry constraints, are inherently reflected in the spectral properties of dynamical systems. Incorporating such physical structure into the neural network architecture may therefore enhance the learning of Koopman spectral information. Future work could explore physics-informed designs that embed prior knowledge directly into the learning process, allowing the network to better align with the system’s underlying dynamics. This integration may improve both the interpretability and accuracy of the extracted Koopman representations.
@@ A.1 Source Code
For reproducibility, the source code will be available at the following link.
@@
$\tfrac{d}{dK}\operatorname{tr}\left(\mathbf{v}_{i}^{*}K^{*}GK\mathbf{v}_{i}\right)=\tfrac{d}{dK}\operatorname{tr}\left((K\mathbf{v}_{i})^{*}G(K\mathbf{v}_{i})\right)$
@@ A.4 Discussion on convergence analysis
    The loss $J(\theta)$ is Lipschitz continuous in $\theta$, and the neural network $\Psi(x;\theta)$ has bounded gradients, facilitating gradient-based optimization.
@@ A.9.1 Dataset details and experimental setup
The dataset utilized in this study is part of the open dataset provided for the ’Sensorium 2023’ competition [39]. The dataset consists of calcium imaging recordings from the primary visual cortex of mice. During the experiments, the mice were presented with natural video stimuli while the activity of thousands of neurons was recorded. The objective of the competition is to predict large-scale neuronal population activity in response to different frames of the stimulus videos, based on the hypothesis that population dynamics in the primary visual cortex, driven by visual stimuli, encode significant information about the dynamics of the videos [3, 29, 15].

### v4

@@ Abstract
Analyzing the long-term behavior of high-dimensional nonlinear dynamical systems remains a significant challenge. While the Koopman operator framework provides a powerful global linearization tool, current methods for approximating its spectral components often face theoretical limitations and depend on predefined dictionaries. Residual Dynamic Mode Decomposition (ResDMD) advanced the field by introducing the spectral residual to assess Koopman operator approximation accuracy; however, its approach of only filtering precomputed spectra prevents the discovery of the operator’s complete spectral information, a limitation known as the ‘spectral inclusion’ problem. We introduce ResKoopNet (Residual-based Koopman-learning Network), a novel method that directly addresses this by explicitly minimizing the spectral residual to compute Koopman eigenpairs. This enables the identification of a more precise and complete Koopman operator spectrum. Using neural networks, our approach provides theoretical guarantees while maintaining computational adaptability. Experiments on a variety of physical and biological systems show that ResKoopNet achieves more accurate spectral approximations than existing methods, particularly for high-dimensional systems and those with continuous spectra, which demonstrates its effectiveness as a tool for analyzing complex dynamical systems.
@@ 1 Introduction
To address these limitations, Residual Dynamic Mode Decomposition (ResDMD) [8] was introduced to offer convergence guarantees through a spectral residual that measures how well the estimated Koopman spectra are. By assessing convergence of spectral residual, ResDMD eliminates spurious spectral components that do not represent the system’s true dynamics, which enhances the reliability of spectral estimation. ResDMD also introduces the concept of pseudospectrum, regions where the spectral content is significant even without discrete eigenvalues, which is particularly valuable for analyzing systems with continuous spectra. However, ResDMD only filters precomputed polluted spectra, which limits its ability to independently refine spectral estimates.
@@ 3.1 ResDMD Review
$=\sum_{i,j=1}^{N_{K}}\bar{v}_{i}[\langle\mathcal{K}\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}-\lambda\langle\psi_{i},\mathcal{K}\psi_{j}\rangle_{\mu}-\bar{\lambda}\langle\mathcal{K}\psi_{i},\psi_{j}\rangle_{\mu}+|\lambda|^{2}\langle\psi_{i},\psi_{j}\rangle_{\mu}]v_{j},$ (3.1)
@@
$\begin{split}\widehat{\textit{res}}(\lambda,\phi)^{2}\coloneqq\frac{1}{m}\mathbf{v}^{*}[\Psi_{Y}^{*}\Psi_{Y}-\lambda(\Psi_{X}^{*}\Psi_{Y})^{*}-\bar{\lambda}\Psi_{X}^{*}\Psi_{Y}+|\lambda|^{2}\Psi_{X}^{*}\Psi_{X}]\mathbf{v}.\end{split}$ (3.3)
@@ 3.2 Residual-based Koopman-learning Network (ResKoopNet)
General framework In this section, we present the Residual-based Koopman-learning Network (ResKoopNet), which minimizes spectral residual to address the spectral inclusion problem and yield a more accurate and complete Koopman spectrum. The overall pipeline is illustrated in Figure 1. ResKoopNet optimizes the dictionary functions by minimizing the total spectral residual $J$ as following
@@
Figure 1: The optimizer on the right searches for the optimal dictionary by minimizing Eq. (3.3).
@@ 4 Application in physical and biological systems
In this section, we present three examples to demonstrate ResKoopNet’s effectiveness in estimating key components of Koopman Mode Decomposition: spectrum, eigenfunctions, and Koopman modes. In the pendulum example, we compare with EDMD, EDMD-DL, Hankel-DMD, and ResDMD, and show that ResKoopNet captures the full Koopman spectrum with significantly fewer dictionary observables than required by ResDMD, while other methods either produce spectral pollution or miss the continuous spectrum. Our turbulence analysis demonstrates ResKoopNet’s ability to detect not only acoustic vibration and turbulent fluctuation (as in ResDMD [8, Section 4.3.1, Section 6.3]) but also to recover the fundamental pressure field structure through the lowest-residual Koopman mode, which is a feature not captured by Hankel-DMD using the same dataset. In the neural dynamics example, we compare against three methods across five mice datasets, with ResKoopNet achieving superior clustering of neural states corresponding to different visual stimuli, as quantified by the Davies-Bouldin index. Note that we include Hankel-DMD [1] as a consistent benchmark across all experiments due to its theoretical guarantees and applicability to both low and high-dimensional systems without requiring predefined dictionaries. It uses time-delayed state measurements by Takens Embedding Theorem [35] instead of using a pre-defined dictionary or applying Galerkin approximation as in EDMD and ResDMD methods (see Appendix A.8.1 for more details).
@@ 4.1 Pendulum
As shown in Figure 2, ResKoopNet requires only $N_{K}=300$ observables to approximate the full spectrum. While the original ResDMD work reported using 964 observables [8, Section 4.3.1], our additional tests indicate that approximately 460 basis functions can produce reasonable circular approximations with ResDMD (see Figure 7 in Appendix A.7.1). Nevertheless, ResKoopNet still achieves comparable accuracy with significantly fewer observables. This efficiency persists even with larger sizes of data points (Figure 3), where $N_{K}=350$ observables remain sufficient. Figure 4 compares ResKoopNet with four classical methods using the 90-initial-point dataset. Both EDMD and ResDMD employ Hermite functions (order 15) for $\theta$ and Fourier functions (order 20) for $\dot{\theta}$, while Hankel-DMD uses a time delay of 150. Though Hankel-DMD produces eigenvalues near the unit circle, it only captures the discrete point spectrum and cannot reconstruct the continuous spectral components. We include Hankel-DMD here for consistency in benchmarking across all experiments, despite its limitations in higher-dimensional systems examined later. The shaded region in the ResDMD plot represents the $\epsilon$-pseudospectrum, which approximates regions where the spectral residual is smaller than a threshold value; however, with 300 basis functions, ResDMD fails to fully recover the unit circle. Overall, this pendulum example demonstrates that ResKoopNet achieves accurate and complete spectral recovery using much fewer observables than other classical methods.
@@ 4.2 Turbulence
Recovering spatial patterns is a typical goal of DMD-based methods, especially in fluid dynamics [32, 27]. Figure 5(a) shows the ground truth pressure distribution, in which a clear asymmetry exists between the upper and lower airfoil surfaces. This fundamental spatial pattern is critical for analyzing aerodynamic performance and structural integrity. While the original Kernel-ResDMD method has successfully detected acoustic vibration and turbulent fluctuation shown in Colbrook and Townsend [8, Section 6.3], it fails to recover the global spatial structure of the pressure field. In contrast, ResKoopNet achieves a significant improvement by accurately reconstructing this dominant spatial pressure field structure itself from its leading Koopman mode as shown in Figure 5(b), which contains essential information about the flow characteristics around the airfoil.
@@ 4.3 Neural dynamics identification in mice visual cortex
Since ResKoopNet directly minimizes the spectral residuals based on eigenfunctions, its estimated evolution of eigenfunctions over time should ideally capture latent dynamics. To evaluate how effectively ResKoopNet reveals latent temporal dynamics in real data, we apply it to a dataset of high-dimensional neural signals and demonstrate its advantages over a series of classical methods: the Hankel-DMD, EDMD (combined with RBF basis) and Kernel-ResDMD. These methods are selected as representative approaches for handling high-dimensional data.
@@
We applied ResKoopNet and three classical Koopman decomposition methods (Hankel-DMD, EDMD with RBF basis, and Kernel-ResDMD) to this dataset, using different implementations and Koopman subspace dimensions. For ResKoopNet, we trained dictionaries on all snapshots from each mouse to avoid overfitting. We reduced the data to 24 dimension via truncated SVD, together with the constant and 25 trainable bases, resulting in 50 eigenfunctions. The neural network we use here consists of 3 hidden layers, each with 200 neurons. The hyperparameter scanning results are included in Appendix Figure 17, which demonstrates that our parameter choice is within the robust range. The decomposed eigenfunctions are shown in Figure 6A(top), with markers indicating ground truth state separations. For Hankel-DMD, we built a Hankel matrix with a delay of 50, producing 50 eigenfunctions per trial. In EDMD with RBF basis, we used the SVD-truncated 300 basis and 1000 RBF functions, resulting in 1301 eigenfunctions. For Kernel-ResDMD, we used normalized Gaussians as kernel functions, setting the Koopman subspace dimension to 299 eigenfunctions based on Colbrook et al. [9]. See Appendix A.9.4 for method details and Appendix A.10 for dictionary size justification. These eigenfunctions, shown in Figure 6A(bottom), Appendix Figure 14A, and Appendix Figure 15A, are compared to ground truth trial identities.
In this study, Koopman eigenfunctions are interpreted as representing dynamical features corresponding to the video stimuli. Consequently, their effectiveness in capturing these key, stimulus-related dynamics is evaluated by assessing how well eigenfunctions from trials with the same video stimulus cluster together and remain distinguishable from those associated with different video stimuli. This effectively transforms the problem into a clustering task based on the separability of eigenfunctions across different stimuli. It is noteworthy that for the ResKoopNet case, averaged trial differences are visibly clear. It is also worth clarifying that clustering these processes is not biologically necessary, as simple clustering of mean firing rates is sufficient (see Appendix Figures 101112). However, this dataset provides an ideal benchmark with clear ground truth labels, enabling accurate comparison of Koopman eigenfunction estimation methods. Effective estimation of Koopman eigenfunctions can be valuable for more complex tasks, such as unsupervised latent state identification or decoding object dynamics from video stimuli, which offers promising directions for future research.
@@ 5 Conclusion and future work
In this paper, we introduced ResKoopNet, a novel approach that minimizes spectral residuals to accurately approximate Koopman operators for complex dynamical systems. This method successfully resolves the spectral inclusion problem while capturing both discrete and continuous spectra with high accuracy. Our experiments across physical and biological systems demonstrate that ResKoopNet significantly outperforms existing methods in identifying coherent structures and uncovering latent dynamics, particularly in high-dimensional applications.
Despite these advantages, ResKoopNet has several limitations. First, the use of neural networks increases computational cost compared to classical numerical algorithms (see Appendix A.6). Second, unlike stochastic approaches such as VAMP [24] and SDMD [46], ResKoopNet currently does not account for stochasticity (see Appendix A.5 for comparison discussion). In addition, its performance can vary depending on the neural architecture and training configuration.
There are several promising directions to address these limitations and extend ResKoopNet’s capabilities. Many physical laws, such as conservation principles or symmetry constraints, are inherently encoded in the spectral properties of dynamical systems. Incorporating such domain knowledge directly into the neural network architecture could significantly enhance the learning of Koopman spectral information. Future work could explore physics-informed neural network designs, develop efficient training strategies to reduce computational costs, and extend the framework to handle stochastic dynamics. Ultimately, ResKoopNet and its future variants have the potential to become powerful tools for understanding and predicting the behavior of complex dynamical systems across diverse scientific and engineering applications, and bridge the gap between data-driven methods and physical principles.
@@ A.1 Source Code
For reproducibility, the source code will be available at this link.
@@
$\tfrac{d}{dK}\operatorname{tr}\left(\mathbf{v}_{i}^{*}K^{*}GK\mathbf{v}_{i}\right)=\tfrac{d}{dK}\operatorname{tr}\left((K\mathbf{v}_{i})^{*}G(K\mathbf{v}_{i})\right).$
@@ A.4 Discussion on convergence analysis
    The loss $J(\theta)$ is strongly convex and Lipschitz continuous in $\theta$, and the neural network $\Psi(x;\theta)$ has bounded gradients, facilitating gradient-based optimization.
@@ A.9.1 Dataset details and experimental setup
The dataset utilized in this study is part of the open dataset provided for the ‘Sensorium 2023’ competition [39]. The dataset consists of calcium imaging recordings from the primary visual cortex of mice. During the experiments, the mice were presented with natural video stimuli while the activity of thousands of neurons was recorded. The objective of the competition is to predict large-scale neuronal population activity in response to different frames of the stimulus videos, based on the hypothesis that population dynamics in the primary visual cortex, driven by visual stimuli, encode significant information about the dynamics of the videos [3, 28, 14].